# Affordance-Guided Reinforcement Learning via Visual Prompting

### v3

@@ Abstract
Robots equipped with reinforcement learning (RL) have the potential to learn a wide range of skills solely from a reward signal. However, obtaining a robust and dense reward signal for general manipulation tasks remains a challenge. Existing learning-based approaches require significant data, such as human demonstrations of success and failure, to learn task-specific reward functions. Recently, there is also a growing adoption of large multi-modal foundation models for robotics that can perform visual reasoning in physical contexts and generate coarse robot motions for manipulation tasks. Motivated by this range of capability, in this work, we present Keypoint-based Affordance Guidance for Improvements (KAGI), a method leveraging rewards shaped by vision-language models (VLMs) for autonomous RL. State-of-the-art VLMs have demonstrated impressive reasoning about affordances through keypoints in zero-shot, and we use these to define dense rewards that guide autonomous robotic learning. On real-world manipulation tasks specified by natural language descriptions, KAGI improves the sample efficiency of autonomous RL and enables successful task completion in 20K online fine-tuning steps. Additionally, we demonstrate the robustness of KAGI to reductions in the number of in-domain demonstrations used for pre-training, reaching similar performance in 35K online fine-tuning steps.222Project page: sites.google.com/view/affordance-guided-rl
@@ I Introduction
Recent advances in large language models (LLMs) and vision-language models (VLMs) trained on Internet-scale data show promising results in using commonsense understanding to plan and reason [1, 2, 3, 4, 5]. They can break down complex instructions provided in natural language into actionable task plans [6, 7, 3, 8, 9, 10], perform visual reasoning in a variety of contexts [11, 12, 13], and even generate coarse robot motions for simple manipulation tasks [7, 14, 15, 9, 16, 13]. However, state-of-the-art models still struggle with understanding 3D physical dynamics and interactions, which is essential to robotic control. Several prior works have utilized large pre-trained models for robotic control, either through few-shot prompting or fine-tuning of large models to generate actions directly [7], plans [3, 10, 9], or code [17, 18]. However, fine-tuning typically requires extensive human supervision, such as teleoperated demonstrations, feedback on whether the task was successfully completed, or a predefined set of skills and their controllers.
An alternative paradigm for fine-tuning robotic policies is autonomous reinforcement learning (RL), which only requires a reward signal to refine the robot’s behavior and can therefore require less supervision in comparison. A significant amount of recent work has also focused on improving the sample efficiency of these algorithms by pre-training on large offline datasets [19, 20, 21, 22, 23, 24]. Despite these advances, obtaining a reward signal is still a non-trivial problem, requiring either careful engineering or large amounts of data to learn a robust reward function [25, 26, 27, 28, 29]. The potential of using pre-trained VLMs to define rewards is hence attractive, but thus far, they have primarily been used for generating sparse rewards [30, 31, 24], which often leads to less efficient learning. VLMs hold much richer and denser knowledge that we can elicit, such as reasoning about the affordances of various objects and environments. In this work, we leverage this understanding to shape rewards for robotic RL.
@@
We present Keypoint-based Affordance Guidance for Improvements (KAGI), a method for open-vocabulary visual prompting to extract rewards from VLMs for autonomous RL. We leverage insights on effective visual prompting techniques [32, 11] to generate affordance keypoints and waypoint trajectories, from which we derive dense shaping rewards for online fine-tuning. As shown in Fig. 1, we integrate our pipeline of extracting affordance keypoint representations from VLMs and computing waypoint-based dense rewards into RoboFuME [24], an autonomous RL system that uses sparse rewards from a fine-tuned VLM. We demonstrate improved success rates on complex object manipulation tasks that the system using only sparse rewards struggles to generalize to, with reduced reliance on in-domain expert demonstrations.
@@ II-A Foundation Models for Robotics
Planning and high-level reasoning. The rapid development of foundation models in recent years has drawn significant attention [33]. This surge of interest has arisen because foundation models demonstrate that models trained on broad, Internet-scale data are highly adaptable to a wide range of downstream tasks. Robotics is a specific downstream task of foundation models that has garnered a lot of interest in the academic community. Foundation models have enabled high-level planning and reasoning for embodied agents [9, 34, 3, 10, 35, 36, 37], generated sub-goals [38, 39, 40], and generated coarse motions for manipulation [41, 9, 13, 14, 15, 16]. These impressive results across a wide range of domains demonstrate great potential in harnessing the rich knowledge and reasoning capabilities of foundation models to enable embodied agents to complete long-horizon tasks.
@@
Despite the rapid development of LLMs for robotic control, recent works have investigated using multi-modal models to jointly provide visual and language prompts to robotic agents [6, 13, 16, 42]. These approaches recognize that converting visual observations into language descriptions and planning solely with language loses rich information associated with the visual modality for scene understanding, a major limitation of using LLMs for spatial planning, reasoning, and task completion. Though LLMs can provide general priors for planning and reasoning, they must be properly grounded to generate accurate responses about real world environments [43, 44, 45, 46, 47]. Our work explores harnessing state-of-the-art VLMs to facilitate reasoning in both language and image domains. We leverage the image domain by computing rewards in image space, using this to bridge high-level LLM plans with low-level actions for embodied tasks.
@@
Spatial reasoning with VLMs. The clear advantage of language is the natural interface for providing task instructions and specifying goals. That said, robotics relies heavily on accurately perceiving and interacting with the environment. Modern VLMs, such as GPT-4V [48] and Gemini [49], demonstrate promising capabilities in combining reasoning with environment perception via visual inputs. Recent works have leveraged this capability to perform keypoint-based reasoning [32, 50] and mark-based visual prompting for zero-shot manipulation [32, 13, 11]. Such approaches facilitate the process of translating high-level plans into low-level robot actions. Beyond zero-shot manipulation, our work explores deriving rewards from VLMs to improve online RL.
@@
Modern VLMs can derive rewards in image space by determining success or failure based on image observations. These reward signals can be used for learning state-action mappings via RL, enabling robots to learn through trial-and-error, without training skill policies via imitation learning which are costly and difficult to scale. VLMs can also guide learning by generating shaping rewards in the form of waypoint trajectories. A key contribution of our work is using VLMs to both determine task completion and specify intermediate waypoints as goals. A major engineering challenge is tuning the inputs to VLMs, which are highly expressive but also opaque, to derive useful reward signals for learning. Both language and image prompts must be tuned to generate accurate and meaningful outputs, as modern VLMs still struggle somewhat with spatial reasoning using raw inputs. Our work explores using the outputs of preliminary VLMs (e.g., bounding boxes [51] and segmentations [52]) to augment raw image observations as guidance for semantic reasoning in more advanced VLMs, thereby leveraging pre-trained representations in VLMs as reward predictors.
@@ II-B Autonomous Reinforcement Learning
Online RL is the paradigm by which an agent gathers data through interaction with the environment, then stores this experience in a replay buffer to update its policy. This contrasts with offline RL, where an agent updates its policy using previously collected data without itself interacting with the environment. A longstanding goal is autonomous RL, an agent that autonomously gathers real-world experience, which holds great promise for scalable robot learning. Agents not only learn through their own experience, but also do not require manual environment resets between trials [53].
Algorithms for autonomous RL are difficult to implement in the real world, with the primary challenges being sample complexity, providing well-shaped rewards, and continual reset-free training. Several works have developed reset-free systems that reduce or eliminate human interventions [54, 24, 55, 56], but reward shaping is an open problem. Manually specified rewards are difficult to engineer and easy to exploit. While hand-designing reward functions is challenging, there is great potential to learn rewards from previously collected data or large pre-trained models. The large bank of offline image and video datasets, plus the high inference speed and accessibility of large pre-trained models, could provide more precise and informative shaping rewards.
RoboFuME [24] is a recent work that fine-tunes MiniGPT-4 [57] to generate sparse task completion rewards for autonomous RL. However, autonomous RL suffers when reward signals are too sparse, and fine-tuning MiniGPT-4 requires a substantial number of in-domain demonstrations per task, which is costly but also makes the system more brittle and less robust to generalization. We draw upon recent work that explores extracting rewards from LLMs [58, 59] and VLMs to guide zero-shot robotic manipulation [32, 31, 30] and online adaptation [24, 60]. In particular, our work leverages affordance representations extracted from VLMs to tackle the dense reward shaping problem.
@@ III-B Keypoint-Based Reward Estimation via Visual Prompting
We take the first observation in the episode , consisting of a top-down image and a side view image . Our goal is to create a candidate set of grasp and target keypoints that GPT-4V can select from. We preprocess these images by (1) passing through GroundingSAM to get segmentations of relevant objects and sample five points from their masks, and (2) overlaying a grid and the sampled keypoints on top of to get . For the side-view image, we augment with evenly-spaced labeled horizontal lines to get , providing depth information not available in the top-down view. See Figures 3(a) and 3(b) as examples. We use this density of grid lines as [32] and preliminary tests found that this was most suitable for GPT-4V to reason with. We pass and with a language instruction and metaprompt to GPT-4V, which generates block_sequence. This sequence is a list of tuples , where is a grid point chosen from representing a position in the xy-plane from top-down and is chosen from representing height in the z-axis from the side view. Dense rewards are calculated with respect to block_sequence in the next step.
For each frame in the episode, we compute the robot position in image space. We use two fitted RANSAC regressors, one for each camera view, to compute the robot position , where is from the top-down and is from the side view. Optionally, we can use pixel trackers [61] to track a specific point on the object to additionally define rewards based on object poses. Using these coordinates, we compute the nearest block to the robot position in block_sequence, (see Figure 3(c) for an example). We compute a reward is based on the L2 distance from the robot position to the block after the closest block in block_sequence, . We use the next block to encourage progression towards the goal and avoid stagnating at the current position. We transform the distance such that reward is between and : . The reward is when the sparse reward is , otherwise it is our dense reward. This reward is then optimized with an online RL algorithm.
We query GPT-4V once at the start of each experiment to generate the affordance keypoints and intermediate waypoints for the forward and backward tasks, which we use for all forward and backward episodes. This is primarily because querying GPT-4V after every episode incurs a prohibitively high cost, but we found that the trajectory is coarse enough to be robust to small randomizations in environment resets.
@@ IV-A Experimental Setup
Tasks. We study a simulated Bin-Sorting task [22, 24], where a WidowX 250 arm needs to sort objects into the correct bin specified by the language instruction. This environment is built in the PyBullet simulation framework. On a physical WidowX 250 robot arm, we study Cloth Covering, Spatula Pick-Place, and Cube Stacking (visualized in Fig. 4). The Cloth Covering task was introduced in RoboFuME [24], and we introduce the latter two tasks to evaluate the applicability of KAGI to new tasks. Each task consists of a forward component (e.g., uncovering the block from under the cloth) and backward component (e.g., covering the block with the cloth), and therefore can be autonomously practiced by alternating between the two sub-tasks.
Comparisons. To understand the performance without any online fine-tuning, we compare KAGI to language-conditioned behavior cloning and a pre-trained RL baseline, CalQL [64]. We also compare to RoboFuME [24], which further trains the pre-trained policy online with a sparse reward derived from a VLM success classifier. This comparison allows us to understand the benefits of the proposed dense rewards. Finally, we evaluate MOKA [32], which executes the trajectory defined by the VLM-generated waypoints. This last comparison does not perform any additional learning.
Datasets. All methods except for MOKA are pre-trained on a combination of trajectories selected from the Bridge dataset [62, 63] and a set of in-domain demonstrations. The in-domain demonstrations consist of forward trajectories and backward trajectories. Following RoboFuME [24], we additionally collect failure trajectories, which are used to train the VLM-based success classifier used by RoboFuME and KAGI. In Section IV-D, we test these methods on a 5 reduction in the quantity of in-domain demonstrations.
Evaluation. In both simulation and real-world evaluations, we roll out each policy for the forward task over trials. For the real-world tasks, success is evaluated as follows. For Cube Covering, the entire cube must be uncovered from the camera perspective shown in Fig. 4 (left); for Spatula Pick-Place, the spatula must be on the plate; for Cube Stacking, the yellow cube must be stable atop the red cube.
@@
Cloth Covering | 35% | 55% | 75% | 75% |
Spatula Pick-Place | 30% | 25% | 40% | 45% |
Cube Stacking | 10% | 15% | 25% | 30% |
@@ IV-B Simulation Experiments
We conduct a preliminary test of our dense reward formulation in a simulated Bin-Sorting task, introduced in RoboFuME [24]. Following their experimental setup, both RoboFuME and KAGI are provided with forward and backward demonstrations, failure demonstrations, and total demonstrations of other pick and place tasks of diverse objects. In simulation, rewards are computed in pixel image space. We evaluate and compare three reward formulations: dense waypoint-based only, sparse success-based only (RoboFuME), and dense + sparse (KAGI).
As shown in Fig. 5, our reward formulation achieves higher success rates than the sparse-only reward formulation that RoboFuME uses. The dense-only reward formulation performs significantly worse than the other two. This indicates that both types of rewards are necessary for efficient and successful learning: while dense rewards are useful in shaping learned behaviors, sparse rewards are also crucial to distinguish truly successful trajectories. We further test our reward formulation succeeds with a reduced number of demonstrations of the task. We test the effect of a reduction in demonstrations ( forward and backward demonstrations) and a reduction in demonstrations ( forward and backward demostrations). In Fig. 5, we see that our reward formulation is robust to these reductions in quantity of task demonstrations.

### v4

@@ Abstract
Robots equipped with reinforcement learning (RL) have the potential to learn a wide range of skills solely from a reward signal. However, obtaining a robust and dense reward signal for general manipulation tasks remains a challenge. Existing learning-based approaches require significant data, such as human demonstrations of success and failure, to learn task-specific reward functions. Recently, there is also a growing adoption of large multi-modal foundation models for robotics that can perform visual reasoning in physical contexts and generate coarse robot motions for manipulation tasks. Motivated by this range of capability, in this work, we present Keypoint-based Affordance Guidance for Improvements (KAGI), a method leveraging rewards shaped by vision-language models (VLMs) for autonomous RL. State-of-the-art VLMs have demonstrated impressive reasoning about affordances through keypoints in zero-shot, and we use these to define dense rewards that guide autonomous robotic learning. On real-world manipulation tasks specified by natural language descriptions, KAGI improves the sample efficiency of autonomous RL and enables successful task completion in 30K online fine-tuning steps. Additionally, we demonstrate the robustness of KAGI to reductions in the number of in-domain demonstrations used for pre-training, reaching similar performance in 45K online fine-tuning steps.222Project page: sites.google.com/view/affordance-guided-rl
@@ I Introduction
Developing generalist robots that can learn and adapt to new tasks autonomously without extensive human efforts has been a longstanding goal in robotics. Typically, this is achieved by collecting a large amount of demonstration data, often in the same domain and manually teleoperated by a human expert, which is costly and limits generalization. Ideally, robotic systems should allow human users to specify reward functions easily and flexibly with high-level instructions. Reinforcement learning (RL) holds great promise of learning new tasks autonomously by optimizing a reward signal. Prior work has sought to improve the sample efficiency of these algorithms by pre-training on large offline datasets [1, 2, 3, 4, 5, 6]. However, learning robust reward functions still requires careful engineering or large amounts of data [7, 8, 9, 10, 11].
@@
Recent advances in large language models (LLMs) and vision-language models (VLMs) trained on Internet-scale data show promising results in breaking down complex language instructions into task plans [12, 13, 14, 15, 16, 17], performing visual reasoning in varied contexts [18, 19, 20], and generating coarse robot motions for simple manipulation tasks [13, 21, 22, 16, 23, 20]. However, fine-tuning large pre-trained models to perform such tasks typically requires extensive human supervision, such as teleoperated demonstrations or feedback on whether the task was successfully completed. A group of prior work uses the generalization capabilities of LLMs/VLMs to solve open sets of tasks through prompt engineering and in-context learning [24, 25, 26], but these works use external motion planners and skill primitives to generate low-level actions, which require significant engineering efforts and are not robust. Another line of work prompts LLMs/VLMs to generate rewards via code to train policies or plan motions to solve the task [27, 28, 29], though existing approaches are only applied to high-level state spaces, navigation, simulated tasks, or coarse manipulation tasks. These approaches can fall short on real-world manipulation tasks that require sophisticated spatial reasoning over visual inputs.
Defining explicit rewards using pre-trained VLMs is hence an attractive approach for learning manipulation tasks. Thus far, VLMs have mainly been used to generate sparse rewards [30, 31, 6], often leading to less efficient learning. VLMs hold much richer geometric and semantic representations that we can elicit, such as reasoning about the affordances of objects and environments. In this work, we leverage insights on effective visual prompting techniques [26, 18] to generate affordance keypoints and waypoint trajectories, from which we directly derive dense rewards to improve the efficiency of online fine-tuning with RL.
@@
We present Keypoint-based Affordance Guidance for Improvements (KAGI), a method for open-vocabulary visual prompting to extract rewards from VLMs for autonomous RL. Our method extracts affordance keypoint representations zero-shot from VLMs and computes waypoint-based dense rewards (Fig. 1), which we integrate into an autonomous RL system that uses sparse rewards from a fine-tuned VLM [6]. KAGI demonstrates improved success rates on complex manipulation tasks while being robust to reduced in-domain expert demonstrations, which systems using only sparse rewards or open-loop primitives struggle to generalize to.
@@ II-A Foundation Models for Robotics
Planning and high-level task reasoning. The rapid development of foundation models has drawn significant attention [32], demonstrating that models trained on Internet-scale data are highly adaptable to a wide range of downstream tasks, including robotics. Foundation models have enabled high-level planning and reasoning for embodied agents [16, 33, 14, 17, 34, 35, 36], generated sub-goals [37, 29, 38], and generated coarse motions for manipulation [39, 16, 20, 21, 22, 23]. These impressive results show great potential in harnessing the rich knowledge and reasoning capabilities of foundation models for tackling robotic manipulation tasks.
@@
Despite the rapid development of LLMs for robotic control, recent works have investigated using multi-modal models to jointly provide visual and language prompts to robotic agents [12, 20, 23, 40]. These approaches recognize that converting visual observations into language descriptions and planning solely with language loses rich information from the visual modality for scene understanding. Though LLMs can provide general priors for spatial planning and reasoning, they must be properly grounded to generate accurate responses about real-world environments [41, 42, 43, 25, 44]. Our work harnesses state-of-the-art VLMs to facilitate reasoning in both language and image domains. We leverage the image domain by computing rewards in image space to bridge high-level LLM plans with low-level actions for embodied tasks.
@@
Spatial reasoning with VLMs. Language provides a highly natural interface for providing task instructions and specifying goals. That said, robotics relies heavily on accurately perceiving and interacting with the environment. Modern VLMs, such as GPT-4o [45] and Gemini [46], demonstrate promising capabilities combining reasoning with environment perception via visual inputs. Recent approaches have explored using VLMs for generating shaping rewards via code [27, 28], reward estimation [47, 29], or preference-based learning [48]. Our work focuses on directly obtaining rewards zero-shot from VLMs by leveraging the rich geometric and semantic representations that VLMs trained on Internet-scale data possess, such that VLM-generated dense rewards can be used effectively in closed-loop autonomous RL systems. Our approach is inspired by works that use generalizable and accurate keypoint-based reasoning of VLMs for zero-shot manipulation [26, 49, 20, 18], translating high-level plans into low-level robot actions. Beyond zero-shot open-loop manipulation, our work presents a novel application of open-vocabulary visual prompting to improve policies pre-trained on diverse data via online RL fine-tuning with VLM-defined dense rewards.
@@
Modern VLMs can derive rewards in image space for learning state-action mappings via RL, enabling robots to learn through trial-and-error without training skill policies via imitation learning which are costly and difficult to scale. A key contribution of our work is using VLMs to determine task completion (sparse rewards) and specify intermediate waypoints as goals (dense rewards), where prior works focus mainly on the former [30, 31, 6]. Our work explores using mark-based visual prompting (e.g., grid lines and object segmentations [50]) to augment raw image observations and guide semantic reasoning in VLMs, leveraging pre-trained representations in VLMs as zero-shot reward predictors.
@@ II-B Autonomous Reinforcement Learning
Online RL is the paradigm by which an agent gathers data through interaction with the environment, then stores this experience in a replay buffer to update its policy. This contrasts with offline RL, where an agent updates its policy using previously collected data without itself interacting with the environment. A longstanding goal is autonomous RL, an agent that autonomously gathers real-world experience online. This approach holds great promise for scalable robot learning as agents learn through their own experience and do not require manual environment resets between trials [51].
Autonomous RL is difficult to implement in the real world, with the primary challenges being sample complexity, providing well-shaped rewards, and continual reset-free training. Several works have developed reset-free systems that reduce human interventions [52, 6, 53, 54], but reward shaping is an open problem. As manually specified rewards are difficult to engineer and easy to exploit, there is great potential to learn rewards from offline data or large pre-trained models. The large bank of image and video datasets, plus the fast inference speed and accessibility of large pre-trained models, could provide more precise and informative shaping rewards.
RoboFuME [6] is a recent work leveraging a reset-free pre-train fine-tune paradigm to train manipulation policies via autonomous RL. RoboFuME pre-trains offline RL policies on diverse offline data [55, 56] and in-domain demonstrations, then fine-tunes a sparse reward classifier [57] for online policy fine-tuning. However, online RL suffers when reward signals are too sparse, and fine-tuning the reward classifier requires a substantial number of in-domain demonstrations per task, which is costly but also makes the system more brittle and less robust to generalization. Our work similarly focuses on the pre-train fine-tune paradigm, specifically on improving the efficiency and robustness of online fine-tuning, enabling effective adaptation of RL policies pre-trained on diverse datasets. To do so, inspired by recent work extracting rewards from VLMs to guide zero-shot robotic manipulation [26, 31, 30] and online adaptation [6, 58], we leverage affordance representations extracted zero-shot from VLMs to tackle the dense reward shaping problem for online RL.
@@ III-B Keypoint-Based Reward Estimation via Visual Prompting
We take the first observation in the episode , consisting of a top-down image and a side view image . Our goal is to create a candidate set of grasp and target keypoints that GPT-4o can select from. We preprocess these images by (1) passing through SAM 2 [50] to get segmentations of relevant objects and sample five points from their masks, and (2) overlaying a grid and the sampled keypoints on top of to get . For the side-view image, we augment with evenly-spaced labeled horizontal lines to get , providing depth information that is not available in the top-down view. See Figures 3(a) and 3(b) for sample processed inputs. We use this density of grid lines as [26] and preliminary tests found that this was most suitable for GPT-4o to reason with. We pass and with a language instruction and metaprompt to GPT-4o, which generates block_sequence. This sequence is a list of tuples , where is a grid point chosen from representing a position in the xy-plane from top-down and is chosen from representing height in the z-axis from the side view. Dense rewards are calculated with respect to block_sequence in the next step.
For each frame in the episode, we compute the robot position in image space. We use two fitted RANSAC regressors, one for each camera view, to compute the robot position , where is from the top-down and is from the side view. Optionally, we can use pixel trackers [59] to track a specific point on the object to additionally define rewards based on object poses. Using these coordinates, we compute the nearest block to the robot position in block_sequence, (see Figure 3(c) for an example). We compute a reward is based on the L2 distance from the robot position to the block after the closest block in block_sequence, . We use the next block to encourage progression towards the goal and avoid stagnating at the current position. We transform the distance such that reward is between and : . The reward is when the sparse reward is , otherwise it is our dense reward. This reward is then optimized using an online RL algorithm [60].
@@ IV-A Experimental Setup
Tasks. In PyBullet simulation, we study a Bin-Sorting task [4, 6], where a WidowX 250 arm needs to sort objects into the correct bin specified by the language instruction. On a physical WidowX 250 robot arm, we study Cloth Covering, Almond Sweeping, Spatula Pick-Place, and Cube Stacking (visualized in Fig. 4). Our tasks span a range of complex skills involving manipulation of deformables, non-prehensile manipulation, functional grasping of objects handles, and precise grasping and placement of cm small cubes designed to challenge our system. Each of our tasks consists of a forward component (e.g., uncovering the block from under the cloth) and backward component (e.g., covering the block with the cloth), and therefore can be autonomously practiced by alternating between these two task components.
Comparisons. To assess performance without any online fine-tuning, we compare KAGI to language-conditioned behavior cloning and a pre-trained RL baseline, CalQL [61]. We also evaluate RoboFuME [6], which fine-tunes the pre-trained RL policy online with sparse rewards derived from a VLM success classifier. This comparison allows us to understand the benefits of the proposed dense rewards. Finally, we evaluate MOKA [26], which executes the trajectory defined by the VLM-generated waypoints. This last comparison does not perform any additional learning or fine-tuning.
Datasets. All methods except for MOKA are pre-trained on a combination of trajectories selected from the Bridge dataset [55, 56] and a set of in-domain demonstrations. The in-domain demonstrations consist of forward trajectories and backward trajectories. Following RoboFuME [6], we additionally collect failure trajectories, which are used to train the VLM-based success classifier used by RoboFuME and KAGI. In Section IV-D, we test these methods on a reduction in the quantity of in-domain demonstrations.
Evaluation. In simulation and real-world evaluations, we roll out each forward task policy for trials. For the real-world tasks, success is evaluated as follows. For Cube Covering, the entire cube must be uncovered from the camera perspective shown in Fig. 4; for Almond Sweeping, all five almonds must reach the right side of the tray; for Spatula Pick-Place, the spatula must be on the plate; for Cube Stacking, the yellow cube must be stable atop the red cube.
@@
| Cloth Covering | 35% | 55% | 80% | 80% |
| Almond Sweeping | 25% | 45% | 70% | 80% |
| Spatula Pick-Place | 30% | 25% | 45% | 65% |
| Cube Stacking | 10% | 15% | 25% | 45% |
@@ IV-B Simulation Experiments
We conduct a preliminary test of our dense reward formulation in a simulated Bin-Sorting task. Following the experimental setup in RoboFuME [6], each approach is provided forward and backward demonstrations, failure demonstrations, and total demonstrations of other pick-place tasks of diverse objects. In simulation, rewards are computed in pixel image space. We compare three reward formulations: dense waypoint-based only, sparse success-based only (RoboFuME), and dense + sparse (KAGI).
As shown in Fig. 5, our reward formulation achieves higher success rates than the sparse-only reward formulation. The dense-only reward formulation performs worse than the other two. This indicates that both types of rewards are necessary for efficient and successful learning: while dense rewards are useful in shaping learned behaviors, sparse rewards are also crucial to distinguish truly successful trajectories. We further test our reward formulation succeeds with a reduced number of demonstrations of the task. We test the effect of a reduction in demonstrations ( forward and backward demonstrations) and a reduction in demonstrations ( forward and backward demostrations). In Fig. 5, we see that our reward formulation is robust to these reductions in quantity of task demonstrations.